OpenStudy (anonymous):

The Party Room at Penny's Pizza rents for an initial fee of $30 and then $5 per hour. Aislyn's bill for her birthday party was $50. For how many hours did she rent the room?

hero (hero):

30 + 5x = 50 5x = 50 - 30 5x = 20 x = 20/5 x = 4 Aislyn rented the room for only 4 hours and got a bill for $50. I hope she paid it.